# Disentangling the development of collective flow in high energy proton proton collisions with a multiphase transport model

### v1

@@ 3.3 Parton escape effects
Figure 6: Two-particle cumulants in the sub-event method after the low multiplicity events non-flow contribution is subtracted(a) and four-particle cumulants (b) for final state partons with $|\eta|<2.4$ and $30.3<p_{T}<$ 3 GeV$/c$ in pp collisions at $\sqrt{s}=13$ TeV versus the number of reference particles in the flow calculations. Gray band shows the normal calculations while magenta band represents the results with random $\phi$ test.
Figure 7: The $p_{\rm T}$ differential $v_{2}$ (a) and $v_{2}^{sub}$ (b) of identified hadrons with random $\phi$ test based on the sub-event method using pseudorapidity gap $|\Delta\eta|>2$ for high multiplicity events with $M>100$ in pp collisions at $\sqrt{s}=13$ TeV.
@@ 4 Summary
Using a multi-phase transport model with PYTHIA8 initial conditions embedded with the sub-nucleon spatial fluctuations using three constituent quarks, we can successfully reproduce the charged hadron productions and the key features of the second order collective flow behavior observed in pp collisions at $\sqrt{s}=13$ TeV. By turning on the parton and hadron evolutions independently, we disentangle how the collective flow develops in different final state evolution stages of pp collisions. We find that the parton evolution stage is necessary for the creation of a sizable two-particle elliptic flow coefficient $v_{2}\{2\}$ and the negative four-particle cumulant $c_{2}\{4\}$. We also find that the coalescence process coupled to the final state parton evolution stage will predict a significant mass splitting feature of the $p_{\rm T}$ differential elliptic flow for identified hadrons. This strong mass ordering effect can be regarded as an important signature to reveal the existence of deconfined parton matter in high multiplicity pp collisions, considering that other approaches without final state parton evolutions usually deliver only small $v_{2}$ splittings.
We find that the collectivity flow in small systems also receives significant contributions from off-equilibrium effects like the parton escape mechanism. We perform a random $\phi$ test to isolate the escape effect in the development of collective flow in high energy pp collisions. We observe that the escape mechanism accounts for a large fraction of the two particle cumulant achieved in the final state. Interestingly, the appearance of a negative $c_{2}\{4\}$ in high multiplicity pp events is also found to be connected to the escape mechanism during the parton cascade. These observations demonstrate the necessity of considering the off-equilibrium effects when extracting the properties of the parton matter created in pp collisions.
@@ Acknowledgements.
We are grateful to Zhenyu Chen, Xinli Zhao, Shusu Shi, Siyu Tang, Chao Zhang, Xiaoming Zhang and Daimei Zhou for helpful discussions. This work is supported by the National Natural Science Foundation of China (Nos. 11905188, 12275103,12061141008), the National Science Foundation under Grant No. 2012947 and 2310021 (Z.-W. L.), the Fundamental Research Funds for the Central Universities China University of Geosciences (Wuhan) with No. G1323523064.

### v2

@@ 4 Summary
Using a multi-phase transport model with PYTHIA8 initial conditions embedded with the sub-nucleon spatial fluctuations using three constituent quarks, we can successfully reproduce the charged hadron productions and the key features of the second order collective flow behavior observed in pp collisions at $\sqrt{s}=13$ TeV. By turning on the parton and hadron evolutions independently, we disentangle how the collective flow develops in different final state evolution stages of pp collisions. We find that the parton evolution stage is necessary for the creation of a sizable two-particle elliptic flow coefficient $v_{2}\{2\}$ and the negative four-particle cumulant $c_{2}\{4\}$. We also find that the coalescence process coupled to the final state parton evolution stage will predict a significant mass splitting feature of the $p_{\rm T}$ differential elliptic flow for identified hadrons. This strong mass ordering effect can be regarded as an important signature to reveal the existence of deconfined parton matter in high multiplicity pp collisions, considering that other approaches without final state parton evolutions usually deliver only small $v_{2}$ splittings. Therefore, investigating the mass ordering feature in experiments with high precision will be essential to distinguish whether this deconfined parton matter is formed in pp collisions ALICE_talk
We find that the collective flow in small systems also receives significant contributions from off-equilibrium effects like the parton escape mechanism. We perform a random $\phi$ test to isolate the escape effect in the development of collective flow in high energy pp collisions. We observe that the escape mechanism accounts for a large fraction of the two particle cumulant achieved in the final state. Interestingly, the appearance of a negative $c_{2}\{4\}$ in high multiplicity pp events is also found to be connected to the escape mechanism during the parton cascade. These observations demonstrate the necessity of considering the off-equilibrium effects when extracting the properties of the parton matter created in pp collisions.
@@ Acknowledgements.
We are grateful to Zhenyu Chen, Xinli Zhao, Shusu Shi, Siyu Tang, Chao Zhang, Xiaoming Zhang and Daimei Zhou for helpful discussions. This work is supported by the National Natural Science Foundation of China (Nos. 11905188, 12275103, 12061141008, 12322508, 12147101), the National Science Foundation under Grant No. 2012947 and 2310021 (Z.-W. L.), the Strategic Priority Research Program of Chinese Academy of Sciences (No. XDB34030000), the STCSM (No. 23590780100), the Fundamental Research Funds for the Central Universities China University of Geosciences (Wuhan) with No. G1323523064.